what would be the force of interaction between 2 charged spheres??

Grade:12

3 Answers

ashish kumar
14 years ago

Charged spheres can be considered as two particle bodies with some charge because the electrical field by a charged sphere at a distance r from its centre is same as the electrical field at a distance r by a charged particle.

So the force of interaction would be = kq2/r , where q = charge on each sphere

                                                                                             r = distance betwen their centres.

dhwani kapoor
14 years ago

no, electric field is not same as that of a charged particle.

 

ashish kumar
14 years ago

Dear Dhwani,

Electrical field of a charged sphere of radius a  and charge q:

At a distance r < a is given by = kqr / a3

                        r > a is given by = kq / r2

                         r = a is given by = kq / a2

Since here we are talking about interaction between 2 charged spheres thatt is kept at a distance r than obviosly this distance is going to be grater than the radius of the sphere i.e. a and hence we will be using the formula to be kq / r2 which is also the electrical field by a particle of charge q at a distance r.

Thank you.